OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (1,5) and (3,6)

OpenStudy (tkhunny):

Find the slope: (5-6)/(1-3) Now what?

hero (hero):

Re-write y = mx + b as y = xm + b, then insert the points as follows: 5 = 1m + b 6 = 3m + b Now, isolate b in both equations: b = 5 - 1m b = 6 - 3m Now set b = b: 5 - 1m = 6 - 3m Place like terms on the same side: 3m - 1m = 6 - 5 Combine like terms: 2m = 1 Isolate m: m = 1/2 Now find b: b = 5 - 1m b = 5 - 1(1/2) b = 5 - 1/2 b = 4 1/2 Therefore the equation of the line is: y = 1/2 x + 4 1/2 or y = 1/2x + 9/2

OpenStudy (tkhunny):

What? Find the Slope: (5-6)/(1-3) = -1/-2 = 1/2 Write the Equation: (y-5) = (1/2)*(x-1) Done.

hero (hero):

Yes, but you didn't provide any kind of explanation whatsoever.

hero (hero):

I could also have "done" the same thing you did, but no one would learn anything.

OpenStudy (tkhunny):

@MathFrustration Have you heard of the Two-Point Form or the Point-Slope Form? They should be in your arsenal.

hero (hero):

Yes, I have heard of all those, however, my method only requires memorizing one formula: y = mx + b

OpenStudy (tkhunny):

Thank you. That would be why I asked the OP. If you have only an hour for a 30 question exam, you're in trouble!

hero (hero):

Not true. You have to understand that I have included a full set of steps with explanation. A student who is familiar with y = mx + b with practice would be able to do this in little to no time at all.

hero (hero):

Meanwhile, your student would be fumbling around trying to remember point-slope form, two -point form, slope form, etc.

OpenStudy (anonymous):

So, if THE problem was wanting me to plot (-5,0) and (0,3) The slope would be 3/5, Which would make it y=3/5x+/- whatever b is.

hero (hero):

If that was the case, then you would already have b, since it represents the y-intercept.

hero (hero):

So you would have the slope and the y-intercept and would be able to write y = mx + b in no time flat

OpenStudy (tkhunny):

Knowing the intercepts, you can just write it down \[\frac{ x }{ -5 } + \frac{ y }{ 3 } = 1\] Please learn various ways to solve as many problem types as you can. With a little practice, you can begin to judge which was is easiest in which situation.

hero (hero):

Plotting is definitely a great way to avoid having to memorize formulas

OpenStudy (anonymous):

Using Hero's method, if it was (-5,0) and (0,3) I can find the slope is 3/5 then do the b=-5(3/5) and find the b= -3

hero (hero):

And yet one's attempt to introduce as many methods as possible surely won't confuse the student beyond clarification.

OpenStudy (anonymous):

So I got that right then.

hero (hero):

Well, actually, if you use my method correctly, you would realize that (0,3) means x = 0, y = 3. In other words, the y-intercept is 3 without doing any algebra.

OpenStudy (anonymous):

Ok.